Physics of CDF

Pekka K. Sinervo University of Toronto

Lecture # 1 Introduction

1. Physics of proton-antiproton scattering

2. General cross section formula

3. Kinematics

4. The Tevatron Collider and the CDF Experiment

Lecture # 2 W/Z Production and Decay

1. Theory of production and decay

2. Detection of W ±’s and Z∘’s

3. σW and W W/W Z production

Lecture # 3 The W Mass

1. Vector boson masses in SM

2. Measurement techniques

3. Latest CDF results

Lecture # 4 Top Production and Decay

1. Standard Model Top Production

2. Top Quark Decay Phenomenology

3. Dilepton and B-tagging

Lecture # 5 More About Top

1. Top Cross Section

2. Top Mass

3. Vtb

4. What Comes Next

TRIUMF Summer Institute July 1995

Introduction: Proton-Antiproton Scattering

The proton is a complex object

• Consists of valence quarks (uud), sea quarks and gluons • each type of parton has momentum distribution function • p beam is really “broad-band” beam of quarks and gluons • p̄ is primarily beam of anti-quarks and gluons

Interested mostly in “hard scattering”

• Characterised by high pT of interaction

Parton Distribution Functions

Fraction of momentum carried by parton given by pdf’s

• Example is EHLQ (Set I) at Q2 = 5 GeV2

• Derived from deep inelastic scattering

• Note large non-valence quark contribution – Largest uncertainty comes from gluon contribution

QCD predicts these PDF’s to evolve with Q^{2}

• Get “softer” as ŝ increases

• Best measurements come from neutrino and DIS

• Raison d’etres for several experiments

Evolution of PDF’s

QCD predicts that f_{q}(x, Q^{2}) will evolve as Q^{2} increases

• Valence quark distributions become softer

• Sea quark and gluon distributions increase in size

• Evolution predicted by the Altarelli-Parisi equations

As example, Duke-Owens I evolution

• Red is x u(x, Q2) and blue is x d(x, Q2)

• Plotted for Q2 = 4 and Q2 = 5000 GeV2

Physics Processes

Rich physics accessible via pp̄ interactions

1. log s interactions dominate cross section • soft non-perturbative hadronic interaction

• results in large number of low pT particles

2. jet production next largest cross section • dominated by two-jet final states

• well described by perturbative QCD (α2s)

3. “light” heavy quark (c and b) production • next to leading order diagrams significant

• Most copious source of b hadrons

4. Top Quark production

5. Electroweak IVB production • Drell-Yan scattering (qq̄ annihilation)

• W W and W Z production

6. Exotic particles • Allows one to reach very high centre-of-mass energy • Can search for new physics at 100 GeV mass scale

– Has to couple to initial state though

Production Rates

Process  σ (nb)  Signature
Min-bias  5 × 107 Anything!
Jets (pT > 10 GeV/c)  3 × 104 Jet with pT > ∼ 10 GeV/c
b (pT > 1 GeV/c)  3 × 104 b → lνlX or b → J/ψX
W, Z  18  W → lνl or Z → e+ e−, μ+ μ−
t t  0.008  Lepton with jets or dileptons

Jet production “easy” to pick out

• Require at least one energetic calorimeter cluster

• Dominated by dijet final states

Heavy quark (b and c) production is copious

• Have to fight light quark backgrounds

Electroweak IVB production “harder”

• Easy to observe decay modes with e± or μ±

– Fully hadronic decays W, Z → qq̄′ overwhelmed by ∗ non-perturbative QCD background

∗ perturbative QCD jet background

Parton Model Cross Section

The Parton Model prediction for the cross section for

pp̄ → X

given by

∑_{q,q′} ∫_{0}^{1} dx_{1} ∫_{0}^{1} dx_{2} f_{1}(x_{1}, Q^{2})f_{2}(x_{2}, Q^{2})σ̂(qq′ → X)(ŝ),

• σ̂(qq′ → X)(ŝ) is parton level cross section – Contains all non-trivial physics – can calculate in SM

• Q2 is 4-momentum of process (∼ M2 X is typical)

• integration over all values of x1 and x2

This is not complete story

• doesn’t include “shattering” of incoming hadrons – generates soft “underlying event” (30-40 particles)

– similar to non-perturbative log (s) interactions

• No good fundamental theoretical description

• Modelled by parametrisations of data

Parton Level Scattering

Parton level cross section σ̂(qq′ → X)(ŝ) is the physics

• Can use SM to calculate for EWK and perturbative QCD

• Worry about large QCD corrections in initial state

Consider Drell-Yan scattering pp̄ → e^{+} e^{−}

σ̂(ŝ) = 3e2q (\frac{4πα2}{3ŝ} )
⇒ σDY = ∑_{q,q′} ∫{1 \atop 0} dx1 ∫{1 \atop 0} dx2 f1(x1, Q2)f2(x2, Q2) [3e2q (\frac{4πα2}{3x1x2s})]
= 3 (\frac{4πα2}{3s} ) ∑_{q=u,d,s} e2q ∫{1 \atop 0} dx1 ∫{1 \atop 0} dx2 \frac{1}{x1x2} f1(x1)f2(x2)

• Should scale as 1/s

• Will fall rapidly with dilepton mass

• Much larger in pp̄ than pp

Drell-Yan Cross Section

Observed Drell-Yan mass distribution from CDF

• dominated by γ coupling at low mass

• Shows clear Z∘ pole

Demonstrates high mass dominated by physics

• Makes collider good place for new heavy particle searches

Transverse Physics is It

Conventional to describe kinematics using

• pT = p sin θ and ET = E sin θ

• pseudorapidity η ≡ − ln tan (θ/2)

η ≃ y ≡ \frac{E + pz}{E − pz} when m ≪ E

• Low pT particle production uniform in η

On collision-to-collision basis

• pz of initial state (qq̄, gg, gq, etc.) is Pbeam(x1 − x2) – Only known statistically

• Final state particles “close” to beam not detectable – Detectors instrumented to η ∼ ±4 → θ ∼ 1.7∘

– Some fraction of “underlying event” escapes detection

Emphasis is on E_{T}

• Detectors are designed to have uniform response in ET • Triggers and data selection based on ET

Tevatron Collider

Tevatron collides 6 × 6 bunchs of p’s and p̄’s

• Antiprotons stored and cooled in accumulator

• Bunchs injected into Main Ring for boost to 120 GeV • Injected into Tevatron, ramped to 900 GeV

Luminosity about 2× design

• Routinely > 2 × 1031 cm−2s−1

• Beam lifetime 12 − 15 hr

• Delivers 3 − 5 pb−1/week

Goals of CDF Experiment

CDF experiment designed to study all high p_{T} physics

• Use p p interactions at √s = 1.8 TeV

• General purpose 5000 ton detector equipped to – Reconstruct charged tracks

– Reconstruct jets using calorimeter

– Detect neutrinos using missing transverse energy – Identify e± and μ± efficiently

Collaboration of ∼ 450 Ph.D.’s taken data since 1987

• 30 nb−1 in 1987

• 4 pb−1 in 1988-89

• 100 pb−1 in 1992-95

Interest is to study all the physics of pp̄ collisions

• Trigger is primarily on large transverse energy events

• Take advantage of large √s and L

• Evolving to study b and c physics as well

The Fermilab Collider Detectors

Two detectors, CDF and D0 / , study Tevatron collisions

• Spectrometer designs are complementary

• Work somewhat in competition

Main detector features:

• Hermetic Calorimetry

– CDF has traditional Pb and Fe calorimeter ∗ Scintillator readout for central region |η| < 1.1

∗ Gas proportional chamber readout for plug/forward – Coverage over |η| < ∼ 4.2

– Good hermiticity for E̸T measurement

– Provides electron ID and jet detection

Resolutions  D0 /  CDF
EM  0.15√E ⊕ 0.004E  0.13√E ⊕ 0.02E
π±, K±  0.50√E  0.80√E
E̸T x,y  1.08 + 0.019 ∑ ET  −0.58 + 0.74√∑ ET

Detectors (continued)

• Muon chambers outside calorimeter – identify charge tracks that pass through calorimeter

• Central Tracking Systems

– CDF has three-tiered system in 1.4 T field ∗ Central Track Chamber (CTC) for 28 < R < 130 cm ∗ Vertex Chambers (VTX) in 10 < R < 23 cm

∗ Silicon Vertex Detector (SVX) in 3 < R < 12 cm ∗ ⇒ σp/p ∼ 0.001p (in GeV/c)

By design, CDF and D0 / are complementary detectors

• D0 / emphasises calorimetery and muon ID

• CDF emphasises tracking and electron ID

• Reality is that both have strengths and weaknesses – CDF Muon system “thin” and fragmented

– D0 / lacks magnetic field – reduces redundancy

Triggering Issues

Use multi-level trigger to tag events in real-time

• Interaction rate of (50 × 10−27 cm 2) ×(1.5 × 1031 cm −2s−1) = 7.5 × 105 Hz

has to be reduced to of order 1-10 Hz

• Most challenging aspect of experiment – Depends critically on all subsystems

– Need efficiency > 75% to be competitive

Triggering for high p_{T} physics is relatively easy

• For example, consider triggering for t t – Tag high PT (> 15 − 20 GeV/c) lepton(s)

– Missing energy requirements (E̸T > 20 − 25 GeV)

– Multijet topologies

– Trigger efficiencies ≥ 90 − 95% in most cases

Bandwith Limitations

Total event rate to tape limited to ∼ 5 Hz

• About 1 Mb/s of data

• Limited by offline processing/data handling

• Sets trigger cross section of < 350 nb

Bandwidth reduction achieved by 3-level trigger

1. look at calorimeter ET and muon stubs • ET > 5 GeV gets down to 1-2 kHz (100 μb)

• Ask for a muon “stub” with pT > 2 GeV/c

2. look at tracks, calorimeter clusters and correlations • e± candidate is charged track and cal cluster • μ± candidate is charged track and μ stub • Correlated cuts gets down to 30 kHz (2 μb)

• Lepton ID gets S/N ∼ 1/10

3. perform most of full offline reconstruction • Identify lepton, jet and neutrino candidates • get down to ∼ 5 Hz (350 nb)

1992-95 Collider Run

Tevatron Collider has been running since 1992

• Run 1A (1992-93) provided ∼ 20 pb−1

– CDF scanned ∼ 1 × 1012 collisions in 19.3 pb−1

∗ Collected 16 × 106 events on tape

• Run 1B (1994-95) has produced another ∼ 85 pb−1 – Run expected to continue till end of 1995

By far the largest and most sensitive dataset

• For example, a top quark with a mass of 160 GeV/c2 – σt t = 8.2 pb ⇒ 400 produced t t events

– D0 / and CDF would each expect to detect ∗ ∼ 6 dilepton events

∗ ∼ 20 lepton + jet events

CDF Physics Suite

Have studied an enormous wealth of physics

1. QCD tests (gg → gg and qq̄ → X) • inclusive jet production and multijet production • differential cross section measurements

2. EWK processes • Drell-Yan (γ, W, Z) production and decay

• Divector Boson production

3. Heavy Quark production • Have produced 3 × 109 bottom quarks already • World’s largest sample of b → J/ψX

• Only place to study top

4. Exotic searches • Most sensitive SUSY searches

• Frontier of large √s physics

• Pushing mass limits of new physics

Will focus on just a few studies

• Chosen for recent interest

• Illustrates techniques well